OpenStudy (anonymous):

Choose the correct simplification of the expression (3x2y3z4)(2x5y2z3).

OpenStudy (unklerhaukus):

is that \[(3x^2y^3z^4)(2x^5y^2z^3)\]?

OpenStudy (anonymous):

yes @UnkleRhaukus

OpenStudy (unklerhaukus):

if you rearrange that expression it can ok like this , \[(3\times 2)(x^2\times x^5)(y^3\times y^2)(z^4\times z^3)\] can you simplify that ?

OpenStudy (unklerhaukus):

remember that \[a^b\times a^c=a^{b+c}\]